# Prym loci of branched double coverings and generalized Andreotti-Mayer loci

### v1

@@ 1. Introduction
Let $C$ be a projective smooth curve of genus $g$ over the complex numbers $\mathbb{C}$, and let $\phi:\tilde{C}\rightarrow{C}$ be a double covering of $C$ branched at $2n$ points. When $n\geq 1$, the norm map $\operatorname{Nm}:\operatorname{Pic}{(\tilde{C})}\rightarrow\operatorname{Pic}{(C)}$ associated with $\phi$ has the connected kernel $P=\operatorname{Ker}{(\operatorname{Nm})}$, which is an abelian subvariety of dimension $d=g+n-1$ in the Jacobian variety $J_{\tilde{C}}=\operatorname{Pic}^{0}{(\tilde{C})}$. Let $\mathcal{L}\in\operatorname{Pic}{(P)}$ be the restriction of the invertible sheaf $\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})$ associated with the theta divisor $\Theta_{\tilde{C}}$ on $J_{\tilde{C}}$. Then the class $[\mathcal{L}]\in{\operatorname{NS}{(P)}}$ determines the polarization on $P$ of type $\Delta=(\underbrace{1,\dots,1}_{n-1},\underbrace{2,\dots,2}_{g})$, and the polarized abelian variety $(P,[\mathcal{L}])$ is called the Prym variety ([11]) of the branched covering $\phi$. From this construction, we obtain the Prym map
@@
where $\mathcal{R}_{g,2n}$ denotes the moduli space of double coverings of curves of genus $g$ branched at $2n$ points, and $\mathcal{A}_{d}^{\Delta}$ denotes the moduli space of $d$-dimensional polarized abelian varieties of type $\Delta$. When $g=0$, the Prym variety $\operatorname{Prym}_{0,2n}(\phi)=(J_{\tilde{C}},[\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})])$ is the hyperelliptic Jacobian, and the Prym map $\operatorname{Prym}_{0,2n}$ is injective by the Torelli theorem. When $g=1$, it is proven in [8] that the Prym map $\operatorname{Prym}_{1,2n}$ is injective for $n\geq 3$. More generally, the following result holds:
@@ 2. $(2)$
If $n\geq 4$ and $d=g+n-1$, then the Prym locus $\mathcal{P}_{g,2n}\subset\mathcal{A}_{d}^{\Delta}$ is an irreducible component of $\mathcal{N}_{d,n-4}^{\Delta}$.
This is a generalization of Theorem 1.2 (2), because $\mathcal{P}_{0,2n}=\mathcal{H}_{d}\subset\mathcal{A}_{d}^{(1,\dots,1)}$ for $d=n-1$. According to the folk conjecture [7, Conjecture 3.15 (2)], it is expected that $\overline{\mathcal{N}_{d,d-3}^{(1,\dots,1)}\setminus\mathcal{N}_{d,d-2}^{(1,\dots,1)}}=\mathcal{H}_{d}$. More generally we expect that $\overline{\mathcal{N}_{d,n-4}^{\Delta}\setminus\mathcal{N}_{d,n-3}^{\Delta}}=\mathcal{P}_{g,2n}$ for $d=g+n-1$, but this remains an open problem.
This paper proceeds as follows: In Section 2, for the Prym variety $(P,[\mathcal{L}])$ of a branched double covering, we explicitly describe the higher base locus $S(P,\mathcal{L})\subset{P}$, and prove that $\mathcal{P}_{g,2n}\subset\mathcal{N}_{d,n-4}^{\Delta}$. In the proof, we provide the equation of the tangent cone $\mathcal{T}_{\Theta,x}\subset{T_{P,x}=T_{P,0}}$ of $\Theta\in|\mathcal{L}|$ at $x\in{S(P,\mathcal{L})}$. In Section 3, we prove that the equations of the tangent cones $\mathcal{T}_{\Theta,x}$ generate a certain subspace in $\operatorname{Sym}^{2}{T_{P,0}}^{\vee}$. In Section 4, following the approach in [1], we introduce a method to bound the dimension of the irreducible component $\mathcal{P}$ of $\mathcal{N}_{d,n-4}^{\Delta}$. When $\mathcal{P}_{g,2n}\subset\mathcal{P}$, by using independent equations of the tangent cones $\mathcal{T}_{\Theta,x}$, we can provide an upper bound for the dimension of $\mathcal{P}$. Combining this with the result from Section 3, we obtain $\dim{\mathcal{P}}\leq\dim{\mathcal{P}_{g,2n}}$, which implies that $\mathcal{P}=\mathcal{P}_{g,2n}$.
@@ 2. Higher base loci and the tangent cone of theta divisors
We assume that $n\geq 1$ and $g\geq 1$. Let $(P,[\mathcal{L}])$ be the Prym variety of $[\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$. Following Mumford’s paper [11], we describe the base locus of the linear system $|\mathcal{L}|$. By [11, p333], we may assume that
$[-1]_{J_{C}}^{*}\mathcal{O}_{J_{C}}(\Theta_{C})=\mathcal{O}_{J_{C}}(\Theta_{C}),\quad[-1]_{J_{\tilde{C}}}^{*}\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})=\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}}),$
@@
Then we have $\pi^{*}\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})={p_{1}^{*}\mathcal{O}_{J_{C}}(2\Theta_{C})\otimes{p_{2}^{*}\mathcal{L}}}$, where $p_{i}$ denotes the $i$-th projection of $J_{C}\times{P}$. Let $s\in{H^{0}(J_{\tilde{C}},\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}}))}$ be a nontrivial section. Since $n\geq 1$, $\psi:J_{C}\rightarrow{J_{\tilde{C}}}$ is injective, and by [11, p335], we have
$\pi^{*}s=\sum_{i=1}^{2^{g}}p_{1}^{*}\alpha_{i}\otimes{p_{2}^{*}\beta_{i}}\in{H^{0}(J_{C}\times{P},p_{1}^{*}\mathcal{O}_{J_{C}}(2\Theta_{C})\otimes{p_{2}^{*}\mathcal{L}})}$
@@
$\operatorname{Bs}{|\mathcal{L}|}=\{\mathcal{F}\in{P}\mid{\psi(J_{C})\subset t_{\mathcal{F}}^{-1}(\Theta_{\tilde{C}})}\}=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}(t_{\phi^{*}\xi}^{-1}(\Theta_{\tilde{C}}))|_{P}.$
More precisely, $H^{0}(P,\mathcal{L})$ is generated by defining sections of the hypersurfaces $(t_{\phi^{*}\xi}^{-1}(\Theta_{\tilde{C}}))|_{P}\subset{P}$ for $\xi\in\operatorname{Pic}^{0}{(C)}$.
For $[\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$, there exists a unique invertible sheaf $\eta_{\phi}\in\operatorname{Pic}^{n}{(C)}$ such that $\Omega_{\tilde{C}}^{1}=\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi})$. Let
@@
$W_{\tilde{C}}=\{\mathcal{F}\in\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}\mid h^{0}(\mathcal{F})\geq 1\}$
@@
$\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})=t_{\mathcal{F}_{0}}^{*}\mathcal{O}_{\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}}(W_{\tilde{C}}),$
@@
$\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\mathcal{F}\otimes\phi^{*}\xi)\geq{h^{0}(\mathcal{G}\otimes\xi)}\geq{i}.$
When $0<2i\leq{n}$, for $E\in{C^{(n-2i)}}$ and $\eta\in\operatorname{Pic}^{n}{(C)}$, we define
@@ 2. $(2)$
    $\mathcal{G}\otimes\xi\in{K_{i}(\phi_{*}D,\eta_{\phi})}$, and $\phi^{*}p\nleq{D}$ for any $p\in{D}$.
@@
$\mathcal{V}_{\phi_{*}D,\mathcal{G}\otimes\xi}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}\simeq{T_{P_{\phi},\mathcal{F}}}.$
@@
First, assume condition (1). If $\phi^{*}p\leq{D}$ for some $p\in{C}$, then $\mathcal{F}\in{B_{i+1}\cap{P_{\phi}}}$, and by Lemma 2.4, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq{i+1}$, which contradicts the assumption that $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=i$. Thus, $\phi^{*}p\nleq{D}$ for any $p\in{C}$. Since $\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D))}\otimes\mathcal{G}^{\otimes 2}=\Omega_{C}^{1}\otimes\eta_{\phi}$, by Lemma 2.6, we obtain an exact sequence
$0\longrightarrow\mathcal{G}\otimes\xi\longrightarrow\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi)\longrightarrow\Omega_{C}^{1}\otimes\mathcal{G}^{\vee}\otimes\xi\longrightarrow 0$
@@
$i=\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\mathcal{F}\otimes\phi^{*}\xi)\geq h^{0}(\mathcal{G}\otimes\xi)\geq{i},$
it follows that $h^{0}(\mathcal{G}\otimes\xi)=i$ and $h^{1}(\mathcal{G}\otimes\xi)=0$. Therefore, $\mathcal{G}\otimes\xi\in{K_{i}(\phi_{*}D,\eta_{\phi})}$, because
$h^{0}(\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\mathcal{O}_{C}(-\phi_{*}D)\otimes(\mathcal{G}\otimes\xi)^{\vee})=h^{0}(\mathcal{G}\otimes\xi^{\vee})$
$=$ $h^{0}(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee}\otimes\xi)+i=h^{0}(\mathcal{F}\otimes\phi^{*}\xi)-h^{0}(\mathcal{G}\otimes\xi)+i=i.$
@@
$\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\mathcal{F}\otimes\phi^{*}\xi)=h^{0}(\mathcal{G}\otimes\xi)=i.$
Let $\sigma:\tilde{C}\rightarrow\tilde{C}$ be the covering involution of $\phi$. The injective homomorphisms
$\mathcal{G}\otimes\xi\hookrightarrow\phi_{*}(\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}(\mathcal{G}\otimes\xi))=\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi)$
@@
$\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\mathcal{O}_{C}(-\phi_{*}D)\otimes(\mathcal{G}\otimes\xi)^{\vee}=\mathcal{G}\otimes\xi^{\vee}\\ \hookrightarrow\phi_{*}(\mathcal{O}_{\tilde{C}}(\sigma^{*}D)\otimes\phi^{*}(\mathcal{G}\otimes\xi^{\vee}))=\phi_{*}(\Omega^{1}_{\tilde{C}}\otimes(\mathcal{F}\otimes\phi^{*}\xi)^{\vee})$
@@
$H^{0}(\mathcal{G}\otimes\xi)\overset{\simeq}{\longrightarrow}H^{0}(\mathcal{F}\otimes\phi^{*}\xi)$
@@
$H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\mathcal{O}_{C}(-\phi_{*}D)\otimes(\mathcal{G}\otimes\xi)^{\vee})\overset{\simeq}{\longrightarrow}H^{0}(\Omega^{1}_{\tilde{C}}\otimes(\mathcal{F}\otimes\phi^{*}\xi)^{\vee}).$
@@
and by the Riemann-Kempf singularity theorem [9, Theorem 2], the tangent cone of $t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})$ at $\mathcal{F}$ is
$(j^{\vee})^{-1}(\mathcal{V}_{\phi_{*}D,\mathcal{G}\otimes\xi})\subsetneq{H^{0}(\Omega_{\tilde{C}}^{1})^{\vee}}\simeq{T_{\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})},\mathcal{F}}},$
where $j^{\vee}:\,H^{0}(\Omega_{\tilde{C}}^{1})^{\vee}\rightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}$ is the dual of the injective homomorphism
$j:\,H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})\longrightarrow H^{0}(\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi}))=H^{0}(\Omega_{\tilde{C}}^{1}).$
Since $T_{P_{\phi},\mathcal{F}}\subset{H^{0}(\Omega_{\tilde{C}}^{1})^{\vee}}$ is isomorphic to $H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}$ via $j^{\vee}$, the tangent cone of $W_{\xi}$ at $\mathcal{F}$ is
$T_{P_{\phi},\mathcal{F}}\cap(j^{\vee})^{-1}(\mathcal{V}_{\phi_{*}D,\mathcal{G}\otimes\xi})\simeq\mathcal{V}_{\phi_{*}D,\mathcal{G}\otimes\xi}\subsetneq H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee},$
@@
When $n<2(i-1)$, we have $\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}=B_{i}\cap{P_{\phi}}=\emptyset$. When $n\geq 2(i-1)$, for $\mathcal{F}\in\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}$, there exist $D\in\tilde{C}^{(n-2(i-1))}$ and $\mathcal{G}\in\operatorname{Pic}^{g-1+(i-1)}{(C)}$ such that
@@
Since $\deg{\mathcal{G}}=g+i-2$, we can find $\xi_{0}\in\operatorname{Pic}^{0}{(C)}$ such that $h^{0}(\mathcal{G}\otimes\xi_{0})=i-1$ and $h^{0}(\mathcal{G}\otimes\xi_{0}^{\vee})=i-1$. Then, we have $\mathcal{G}\otimes\xi_{0}\in{K_{i-1}(\phi_{*}D,\eta_{\phi})}$. If $\mathcal{F}\notin{B_{i}}$, then $\phi^{*}p\nleq{D}$ for any $p\in{D}$, and by Lemma 2.5, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi_{0}}}=i-1$. This is a contradiction to $\mathcal{F}\in\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}{W_{\xi,\geq{i}}}$, hence $\mathcal{F}\in{B_{i}}$. ∎
Lemma 2.7.
$\dim{(B_{i}\cap{P_{\phi}})}=n-2i$ for $0<2i\leq{n}$. If $0<2i<n$, then $B_{i}\cap{P_{\phi}}$ is irreducible. If $0<2i=n$, then #$\#{(B_{i}\cap{P_{\phi}})}=4^{g}$.
@@
is a finite étale covering of degree $4^{g}$. When $n=2i$,
@@
is bijective, hence #$\#{(B_{i}\cap{P_{\phi}})}=\#{Y_{i}}=4^{g}$. Assume $n\geq 2i+1$. Let $\gamma:[0,1]\rightarrow\tilde{C}$ be a loop on $\tilde{C}$, and $\iota$ be the morphism defined by
@@
Then, we have a unique lift $\delta:[0,1]\rightarrow{H^{0}(\Omega_{C}^{1})^{\vee}}$ of $\iota\circ\phi\circ\gamma$ satisfying $\delta(0)=0$:
@@
$[0,1]\longrightarrow{Y_{i}};\,t\longmapsto(\gamma(t)+D^{\prime},\mathcal{G}\otimes\xi(t)^{\vee})$
@@
$[0,1]\longrightarrow\tilde{C}^{(n-2i)};\,t\longmapsto\gamma(t)+D^{\prime}$

### v2

@@ 1. Introduction
[…]
@@
where $\mathcal{R}_{g,2n}$ denotes the moduli space of double coverings of curves of genus $g$ branched at $2n$ points, and $\mathcal{A}_{d}^{\Delta}$ denotes the moduli space of $d$-dimensional polarized abelian varieties of type $\Delta$. When $g=0$, the Prym variety $)\operatorname{Prym}_{0,2n}(\phi)=(J_{\tilde{C}},[\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}}$)]) is the hyperelliptic Jacobian, and the Prym map $\operatorname{Prym}_{0,2n}$ is injective by the Torelli theorem. When $g=1$, it is proven in [8] that the Prym map $\operatorname{Prym}_{1,2n}$ is injective for $n\geq 3$. More generally, the following result holds:
@@ 2. $(2)$
If $n\geq 4$ and d=g+n1d=g+n-1, then the Prym locus $\mathcal{P}_{g,2n}\subset\mathcal{A}_{d}^{\Delta}$ is an irreducible component of $\mathcal{N}_{d,n-4}^{\Delta}$.
This is a generalization of Theorem 1.2 (2), because $\mathcal{P}_{0,2n}=\mathcal{H}_{d}\subset\mathcal{A}_{d}^{(1,\dots,1)}$ for d=n1d=n-1. According to the folk conjecture [7, Conjecture 3.15 (2)], it is expected that $\overline{\mathcal{N}_{d,d-3}^{(1,\dots,1)}\setminus\mathcal{N}_{d,d-2}^{(1,\dots,1)}}=\mathcal{H}_{d}$. More generally we expect that $\overline{\mathcal{N}_{d,n-4}^{\Delta}\setminus\mathcal{N}_{d,n-3}^{\Delta}}=\mathcal{P}_{g,2n}$ for d=g+n1d=g+n-1, but this remains an open problem.
This paper proceeds as follows: In Section 2, for the Prym variety $(P,[\mathcal{L}])$ of a branched double covering, we explicitly describe the higher base locus $S(P,\mathcal{L})\subset{P}$, and prove that $\mathcal{P}_{g,2n}\subset\mathcal{N}_{d,n-4}^{\Delta}$. In the proof, we provide the equation of the tangent cone $\mathcal{T}_{\Theta,x}\subset{T_{P,x}=T_{P,0}}$ of $\Theta\in|\mathcal{L}|$ at $x\in{S(P,\mathcal{L})}$. In Section 3, we prove that the equations of the tangent cones $\mathcal{T}_{\Theta,x}$ generate a certain subspace in $\operatorname{Sym}^{2}{T_{P,0}}^{\vee}$. In Section 4, following the approach in [1], we introduce a method to bound the dimension of an irreducible component $\mathcal{P}$ of $\mathcal{N}_{d,n-4}^{\Delta}$. When $\mathcal{P}_{g,2n}\subset\mathcal{P}$, by using independent equations of the tangent cones $\mathcal{T}_{\Theta,x}$, we can provide an upper bound for the dimension of $\mathcal{P}$. Combining this with the result from Section 3, we obtain $\dim{\mathcal{P}}\leq\dim{\mathcal{P}_{g,2n}}$, which implies that $\mathcal{P}=\mathcal{P}_{g,2n}$.
@@ 2. Higher base loci and the tangent cone of theta divisors
We assume that $n\geq 1$ and $g\geq 1$. Let $(P,[\mathcal{L}])$ be the Prym variety of $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$. Following Mumford’s paper [11], we describe the base locus of the linear system $|\mathcal{L}|$. By [11, p333], we may assume that
[…]
@@
Then we have $\pi^{*}\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}})={\mathrm{pr}_{1}^{*}\mathcal{O}_{J_{C}}(2\Theta_{C})\otimes{\mathrm{pr}_{2}^{*}\mathcal{L}}}$, where $\mathrm{pr}_{i}$ denotes the $i$-th projection of $J_{C}\times{P}$. Let $s\in{H^{0}(J_{\tilde{C}},\mathcal{O}_{J_{\tilde{C}}}(\Theta_{\tilde{C}}))}$ be a nontrivial section. Since $n\geq 1$, $\psi:J_{C}\rightarrow{J_{\tilde{C}}}$ is injective, and by [11, p335], we have
$\pi^{*}s=\sum_{i=1}^{2^{g}}\mathrm{pr}_{1}^{*}\alpha_{i}\otimes{\mathrm{pr}_{2}^{*}\beta_{i}}\in{H^{0}(J_{C}\times{P},\mathrm{pr}_{1}^{*}\mathcal{O}_{J_{C}}(2\Theta_{C})\otimes{\mathrm{pr}_{2}^{*}\mathcal{L}})}$
@@
$\operatorname{Bs}{|\mathcal{L}|}=\{\mathcal{F}\in{P}\mid{\psi(J_{C})\subset t_{\mathcal{F}}^{-1}(\Theta_{\tilde{C}})}\}=\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}t_{\phi^{*}\xi}^{-1}(\Theta_{\tilde{C}})|_{P}.$
More precisely, $H^{0}(P,\mathcal{L})$ is generated by defining sections of the hypersurfaces $t_{\phi^{*}\xi}^{-1}(\Theta_{\tilde{C}})|_{P}\subset{P}$ for $\xi\in\operatorname{Pic}^{0}{(C)}$.
For $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$, there exists a unique invertible sheaf $\eta_{\phi}\in\operatorname{Pic}^{n}{(C)}$ such that $\Omega_{\tilde{C}}^{1}=\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi})$. Let
@@
$W_{\tilde{C}}=\{\mathcal{F}\in\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})}\mid h^{0}(\tilde{C},\mathcal{F})\geq 1\}$
@@
[…]
@@
$\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)\geq{h^{0}(C,\mathcal{G}\otimes\xi)}\geq{i}.$
The following Mumford’s exact sequence is essential for the proof of Proposition 2.3.
Lemma 2.5 ([11] p338).
of locally free sheaves on $C$.
In the following, we write $H^{i}(\mathcal{G})=H^{i}(C,\mathcal{G})$ for any sheaf $\mathcal{G}$ on $C$. When $0<2i\leq{n}$, for $E\in{C^{(n-2i)}}$ and $\eta\in\operatorname{Pic}^{n}{(C)}$, we define
@@ 2. $(2)$
    $\mathcal{G}\otimes\xi\in{K_{i}(\phi(D),\eta_{\phi})}$, and $\phi^{*}p\nleq{D}$ for any $p\in{C}$.
@@
$\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}\simeq{T_{P_{\phi},\mathcal{F}}}.$
@@
First, assume condition (1). If $\phi^{*}p\leq{D}$ for some $p\in{C}$, then $\mathcal{F}\in{B_{i+1}\cap{P_{\phi}}}$, and by Lemma 2.4, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq{i+1}$, which contradicts the assumption that $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=i$. Thus, $\phi^{*}p\nleq{D}$ for any $p\in{C}$. Since $\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(D))}\otimes\mathcal{G}^{\otimes 2}=\Omega_{C}^{1}\otimes\eta_{\phi}$, by Lemma 2.5, we obtain the exact sequence
$0\longrightarrow\mathcal{G}\otimes\xi\longrightarrow\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi)\longrightarrow\Omega_{C}^{1}\otimes\mathcal{G}^{\vee}\otimes\xi\longrightarrow 0.$
@@
$i=\operatorname{mult}_{\mathcal{F}}{W_{\xi}}\geq\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)\geq h^{0}(\mathcal{G}\otimes\xi)\geq{i},$
it follows that $h^{0}(\mathcal{G}\otimes\xi)=i$ and $h^{1}(\mathcal{G}\otimes\xi)=0$. Therefore, $\mathcal{G}\otimes\xi\in{K_{i}(\phi(D),\eta_{\phi})}$, because
$h^{0}(\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(-D))}\otimes(\mathcal{G}\otimes\xi)^{\vee})=h^{0}(\mathcal{G}\otimes\xi^{\vee})$
$=$ $h^{0}(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee}\otimes\xi)+i=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)-h^{0}(\mathcal{G}\otimes\xi)+i=i.$
@@
$\operatorname{mult}_{\mathcal{F}}{t_{\phi^{*}\xi}^{-1}(W_{\tilde{C}})}=h^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)=h^{0}(\mathcal{G}\otimes\xi)=i.$
Let $\sigma:\tilde{C}\rightarrow\tilde{C}$ be the covering involution of $\phi$. Since $\mathcal{F}\in{P_{\phi}}$, we have
$\Omega_{\tilde{C}}^{1}=\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi})=\mathcal{F}\otimes\sigma^{*}\mathcal{F}=\mathcal{F}\otimes\mathcal{O}_{\tilde{C}}(\sigma^{*}D)\otimes\phi^{*}\mathcal{G}.$
The injective homomorphisms
$\mathcal{G}\otimes\xi\hookrightarrow\phi_{*}(\mathcal{O}_{\tilde{C}}(D))\otimes\mathcal{G}\otimes\xi=\phi_{*}(\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}(\mathcal{G}\otimes\xi))=\phi_{*}(\mathcal{F}\otimes\phi^{*}\xi)$
@@
$\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(-D))}\otimes(\mathcal{G}\otimes\xi)^{\vee}=\mathcal{G}\otimes\xi^{\vee}\\ \hookrightarrow\phi_{*}(\mathcal{O}_{\tilde{C}}(\sigma^{*}D))\otimes\mathcal{G}\otimes\xi^{\vee}=\phi_{*}(\mathcal{O}_{\tilde{C}}(\sigma^{*}D)\otimes\phi^{*}(\mathcal{G}\otimes\xi^{\vee}))=\phi_{*}(\Omega^{1}_{\tilde{C}}\otimes(\mathcal{F}\otimes\phi^{*}\xi)^{\vee})$
@@
$H^{0}(C,\mathcal{G}\otimes\xi)\overset{\simeq}{\longrightarrow}H^{0}(\tilde{C},\mathcal{F}\otimes\phi^{*}\xi)$
@@
$H^{0}(C,\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\operatorname{Nm}{(\mathcal{O}_{\tilde{C}}(-D))}\otimes(\mathcal{G}\otimes\xi)^{\vee})\overset{\simeq}{\longrightarrow}H^{0}(\tilde{C},\Omega^{1}_{\tilde{C}}\otimes(\mathcal{F}\otimes\phi^{*}\xi)^{\vee}).$
@@
[…]
$(j^{\vee})^{-1}(\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi})\subsetneq{H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1})^{\vee}}\simeq{T_{\operatorname{Pic}^{\tilde{g}-1}{(\tilde{C})},\mathcal{F}}}$,
where $j^{\vee}:\,H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1})^{\vee}\rightarrow{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}$ is the dual of the injective homomorphism
$j:\,H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})\longrightarrow H^{0}(\tilde{C},\phi^{*}(\Omega_{C}^{1}\otimes\eta_{\phi}))=H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1}).$
Since $T_{P_{\phi},\mathcal{F}}\subset{H^{0}(\tilde{C},\Omega_{\tilde{C}}^{1})^{\vee}}$ is isomorphic to $H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}$ via $j^{\vee}$, the tangent cone of $W_{\xi}$ at $\mathcal{F}$ is
$T_{P_{\phi},\mathcal{F}}\cap(j^{\vee})^{-1}(\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi})\simeq\mathcal{V}_{\phi(D),\mathcal{G}\otimes\xi}\subsetneq H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee},$
@@
When n<2(i1)n<2(i-1), we have defined that $B_{i-1}=B_{i}=\emptyset$, and it follows that $\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}=\emptyset$ and $B_{i}\cap{P_{\phi}}=\emptyset$. When $n\geq 2(i-1)$, for $\mathcal{F}\in\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}W_{\xi,\geq{i}}$, there exist $D\in\tilde{C}^{(n-2(i-1))}$ and $\mathcal{G}\in\operatorname{Pic}^{g-1+(i-1)}{(C)}$ such that
@@
Since $\deg{\mathcal{G}}=g+i-2$, we can find $\xi_{0}\in\operatorname{Pic}^{0}{(C)}$ such that $h^{0}(\mathcal{G}\otimes\xi_{0})=i-1$ and $h^{0}(\mathcal{G}\otimes\xi_{0}^{\vee})=i-1$. Then, we have $\mathcal{G}\otimes\xi_{0}\in{K_{i-1}(\phi(D),\eta_{\phi})}$. If $\mathcal{F}\notin{B_{i}}$, then $\phi^{*}p\nleq{D}$ for any $p\in{D}$, and by Lemma 2.6, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi_{0}}}=i-1$. This is a contradiction to $\mathcal{F}\in\bigcap_{\xi\in\operatorname{Pic}^{0}{(C)}}{W_{\xi,\geq{i}}}$, hence $\mathcal{F}\in{B_{i}}$. ∎
Proposition 2.7.
$\dim{(B_{i}\cap{P_{\phi}})}=n-2i$ for $0<2i\leq{n}$. If $0<2i<n$, then $B_{i}\cap{P_{\phi}}$ is irreducible. If $0<2i=n$, then ${(B_{i}\cap{P_{\phi}})}=4^{g}$.
@@
is a finite étale covering of degree $4^{g}$, hence $Y_{i}$ is smooth. When $n=2i$,
@@
is bijective, hence ${(B_{i}\cap{P_{\phi}})}=\#{Y_{i}}=4^{g}$. Assume $n\geq 2i+1$. Let $\gamma:[0,1]\rightarrow\tilde{C}$ be a loop on $\tilde{C}$, and let $\iota$ be the morphism defined by
@@
Then, there is a unique lift $\delta:[0,1]\rightarrow{H^{0}(\Omega_{C}^{1})^{\vee}}$ of $\iota\circ\phi\circ\gamma$ such that $\delta(0)=0$:
@@
$0,1]\longrightarrow{Y_{i}};\,t\longmapsto(\gamma(t)+D^{\prime},\mathcal{G}\otimes\xi(t)^{\vee})$
@@
$0,1]\longrightarrow\tilde{C}^{(n-2i)};\,t\longmapsto\gamma(t)+D^{\prime}$